Given a normal distribution with µ = 100 and σ = 10, if you select a sample of n = 25, what is the probability that X bar is
a. less than 95?
b. Between 95 amd 97.5?
c. Above 102.2?
d. There is a 65% chance that X bar is above what value?

Found 2 solutions by stanbon, sbepk4:
Answer by stanbon(75887)

Given a normal distribution with µ = 100 and sigma = 10, if you select a sample of n = 25, what is the probability that X bar is
a. less than 95?
z(95) = (95-100)/10/sqrt25) = -5/2 = -2.5
P(xbar<95)=P(z<-2.5)=0.006
-------------
b. Between 95 amd 97.5
z(95)=-2.5 ; z(97.5) = (97.5-100)/10/5=-2.5/2=-1.25
P(95 --------------
c. Above 102.2?
z(102.2)=(102.2-100)/2=1.1
P(xbar>102.2)=P(z>1.1)=0.135666...
---------------
d. There is a 65% chance that X bar is above what value?
The corresponding z-value is z=-0.38532...
Converting to raw score: -0.38532 = (xbar-100)/2;
xbar=2*(-0.38532)+100
xbar=99.23
-------------
======================
Cheers,
Stan H.

Answer by sbepk4(1)